# Underactuated Control of Multiple Soft Pneumatic Actuators via Stable Inversion

### v1

@@ IV-E Stable Inversion Control Tests
The stable inversion controller designed in Sec. III-C is implemented to control and synchronize the two fingers in this subsection. The result is demonstrated in Fig. 7 (b). The motion of two soft fingers is synchronized and they reach a steady state at nearly the same time with nearly no errors. Compared to open-loop and LQR control, the controller enables more precise and synchronized motions for both soft fingers with a single syringe pump. Their motions are synchronized with around 2 deg differences. Our previous work [7] utilized a syringe pump for each soft finger of the two-finger gripper to reach synchronization. However, one more syringe pump is needed compared to this research. If there are three-finger or four-finger grippers, more syringe pumps are required, which implies more costs, space, and weight. That makes the applications of the soft gripper setup more difficult.
@@ IV-F Disturbance Tests
According to the Theorem 3, the feedback loop is designed to deal with the model errors or disturbances caused by external forces. The disturbance is given at around t = 1.4 sec when the fingers arrive at the steady state as Fig. 7 (c). The external force is only applied to the left finger. The algebraic controller is able to adjust the systems to the desired reference when the external force is applied. The experimental results support the Theorem 3.
@@ V-A Discussion
The proposed control algorithms successfully synchronize the motions of two soft fingers within a soft gripper. The performance is validated by the real-world experimentation. However, there is a limitation of using this control algorithm in which the desired output function of (28) should be in the image space of the model matrix $P(s)$. If the desired output function is out of the image space of $P(s)$, the proposed method is no longer valid. For instance, if the desired function of one finger is $\pi/4$ and another one is $-\pi/4$, the solution of (28) does not exist. The negative bending angle is out of the image space of this soft gripper. Specifically, since the desired functions are in the image space of the model matrix, the position and its derivative are constrained by the output functions. Unlike the linear quadratic regulator, the position and velocity can be adjusted by selecting a suitable weighting matrix function. In short, the synchronization of the desired systems is achievable but the output function cannot be randomly selected.
The existence of the synchronization subspace is discussed in Sec. III-C. The results of the Sec. IV-E and Fig. 7 (b) support the discussion and Definition 2. The open-loop control has a slower input command and the responses of both fingers differ as Fig. 7 (a). The stable inversion control has a faster input command and the motions of the two soft fingers are synchronized as Fig. 7 (b). It, therefore, can be concluded that the (28) is controllable and the synchronization subspace exists. Note that the LQR control also has faster responses but its command trajectory is different from that of stable inversion control. The result of LQR control displays that it does not belong to the synchronization subspace.
@@ V-B Conclusion
This paper explores the underactuated control of multiple fingers within a soft gripper, validating a controller that contain feedforward and feedback loops designed via algebraic methods. The soft fingers are designed based on an optimal design framework and their dynamical models are obtained by applying nonlinear mechanics. The feedforward controller is designed based on stable inversion of the system model matrix, while the feedback loop is incorporated to handle the system perturbations. Simulation results demonstrate the efficacy of the control algorithms in synchronizing both two-finger and three-finger grippers. Experimental validation further confirms the feasibility of synchronizing motions within a two-finger gripper setup. Even if there is a disturbance, the controller is able to adjust the systems to the reference. The control strategy reduces the number of inputs (air pumps) which may benefit the implementation of multi-finger soft grippers.

### v2

@@ IV-E Underactuated Control Tests
Due to the proposed controller, system response is quicker and reduces the uncertainty band of soft actuators. Our previous work [7] utilized a syringe pump for each soft finger of the two-finger gripper to reach synchronization. However, one more syringe pump is needed compared to this research. If there are three-finger or four-finger grippers, more syringe pumps are required, which implies more costs, space, and weight. That makes the applications of the soft gripper setup more difficult.
@@ IV-F Disturbance Tests
According to the Theorem 3, the feedback loop is designed to deal with the model errors or disturbances caused by external forces based on the Theorem 3 ((26) and (28)). The disturbance is generated by a human finger and is given at around t = 1.4 $sec$ when the fingers arrive at the steady state as Fig. 7 (c). The external force is only applied to the left finger. The proposed control algorithm is able to regulate the systems to the desired reference when the external force is applied. The experimental results support the Theorem 3.
@@ V-A Discussion
The proposed control algorithms successfully coordinate the motions of two soft fingers within a soft gripper. The performance is validated by real-world experimentation. However, if systems have identical models (i.e., $P_{1}=P_{2}=...=P_{n_{y}}$ in (6)), those systems will automatically coordinate their motions given an input $U(s)$. Since the models of soft fingers are different (i.e., $P_{1}\neq P_{2}\neq...\neq P_{n_{y}}$ in (6)), this approach is proposed to address this issue and coordinate their motions. Even if the soft fingers have the same dimensional parameters, their system models are different due to the properties of soft materials [16].
Furthermore, there is a limitation due to the hardware configuration of the soft gripper. The soft gripper has a parallel nature of multiple fingers driven by a single syringe pump. If the desired output function is out of the image space of $P(s)$, the soft gripper may not reach the desired states. For instance, if the desired function of one finger is $\pi/4~{}rad$ and another one is $-\pi/4~{}rad$, the solution of the system (24) does not exist. The negative bending angle is out of the image space of this soft gripper.
The proposed approach sets the input-output relation characterized by $H(s)$, either a low-pass filter or designed by loop-shaping. The system responses are influenced by the $H(s)$. By enlarging the bandwidth of $H(s)$, the output moves toward the steady-state angle faster. Thus, the system responses are optimized by designing a suitable $H(s)$ as shown in both Sec.IV-B and Sec. IV-E. The simulation and experimental visualizations are depicted in Fig. 6 (a) and Fig. 7 (b). The steady-state angle of each finger comes closer to the desired angle, and as a result, better coordination is achieved.
@@ V-B Conclusion
This paper explores the underactuated control of multiple fingers within a soft gripper, validating a controller that contains feedforward and feedback loops designed via stable model inversion. The soft fingers are designed based on an optimal design framework and their dynamical models are obtained by applying nonlinear mechanics. The feedforward controller is designed based on stable inversion of the system model matrix, while the feedback loop is incorporated to handle the system perturbations. Simulation results demonstrate the efficacy of the control algorithms in controlling a two-finger gripper. Experimental validation further confirms the feasibility of coordinating motions within a two-finger gripper setup, achieving high-speed transient responses and minimal steady-state errors. Even if there is a disturbance, the controller is able to regulate the systems to the reference. The control strategy reduces the number of inputs (air pumps) which may benefit the implementation of multi-finger soft grippers.